OpenStudy (anonymous):

A line passes through (4, 5) and (8, 9). Which equation best represents the line? y = x + 1 y = 1 y = 2x + 1 y = 5x + 8

OpenStudy (anonymous):

could you help

OpenStudy (arthur_ser):

The general equation of a line is: y=ax + b The points the line passes through must evaluate its equation thus: 5=a*4 + b and 9=a*8 + b Solve the system by substracting each sides: 5-9 =4*a -8*a +b -b -4=-4*a a=1 Substitute a into one the first equations to find b: 5=1*4+ b b=1 so the equation of the line is y=x+1

OpenStudy (anonymous):

thank you